OpenStudy (anonymous):

What is the equation of the line that passes through (–2, –3) and is perpendicular to 2x – 3y = 6?

OpenStudy (texaschic101):

first we need to find the slope of 2x - 3y = 6. The reason we have to do this is because perpendicular lines have negative reciprocal slopes. Do you know how to find the slope of that line ?

OpenStudy (anonymous):

I think the slope is 2? right?

OpenStudy (texaschic101):

2x - 3y = 6 -- subtract 2x from both sides -3y = -2x + 6 -- divide by -3 y = 2/3x - 2 In y = mx + b form, the slope is in the m position...therefore the slope is 2/3. However, we need the negative reciprocal slope. All that means is " flip " the slope and change the sign. So the slope we need, the negative reciprocal, is -3/2. Do you understand so far ?

OpenStudy (texaschic101):

do you have any questions or shall I continue ?

OpenStudy (anonymous):

Continue please

OpenStudy (texaschic101):

now we will use y = mx + b slope(m) = -3/2 (-2,-3)...x = -2 and y = -3 we already found the slope we will be using, so all we need to do is find the y intercept (b) now sub and find b, the y intercept y = mx + b -3 = -3/2(-2) + b -3 = 3 + b --- subtract b from both sides -3 - 3 = b -6 = b so our perpendicular equation is : y = -3/2x - 6 any questions ?

OpenStudy (anonymous):

no thanks bro

OpenStudy (texaschic101):

oops..typo...I meant subtract 3 from both sides...not b...lol